OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1 A. -69x + 10y = 51 B. 69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51

OpenStudy (nurali):

\[y = -6.9x + 5.1\] \[y =\frac{ -69x }{ 10 }+\frac{ 51 }{ 10 }\] \[y=\frac{ -69x+51 }{ 10 }\] \[10y=-69x+51\] \[69x+10y=51\]

OpenStudy (anonymous):

thank youu

OpenStudy (nurali):

My Pleasure.